### v1

@@ 4.4 Experimental Results
[thomas2018predicting] 0.7803 0.8764 0.8008
[abedi2023affect] 0.7637 0.8710 0.7980
[chen2019faceengage] 0.7475 0.8890 0.7973
[abedi2023bag] 0.8191 0.8926 0.9018
@@
5 Conclusion and Future Works
@@
Our research led to the development of a novel deep-learning framework for student engagement measurement. In the proposed framework, sequences of facial landmarks are extracted from consecutive video frames and analyzed by ordinal ST-GCNs to make inferences regarding the engagement level of the subject in the video. The successful application of our model to the EngageNet [engagenet] and Online SE [thomas2022automatic] datasets not only confirms its efficacy but also establishes a new standard in engagement level classification accuracy, outperforming existing methods. As the sole input information to the developed engagement measurement models, the 3D facial landmarks contain information about head pose [malek2021head], AUs [jacob2021facial], eye gaze [grishchenko2020attention], and affect [liu2022graph, toisoul2021estimation], which are the key indicators of behavioral and affective engagement. The relatively lightweight ST-GCN and its integration with real-time MediaPipe make the proposed framework capable of being deployed on virtual learning platforms and measuring engagement in real time. Despite these advancements, the proposed method is not without limitations. The interpretability and explainability of the proposed method could be improved through techniques such as Gradient-based Class Activation Maps [zheng2023skeleton] for ST-GCNs [9746621]. This would facilitate an understanding of which facial landmarks and specific timestamps contribute to certain levels of engagement. To improve the performance of the proposed method, the following direction could be investigated: analyzing facial landmarks with more advanced ST-GCN, which are equipped with attention mechanism [lin2023actionlet], applying augmentation techniques to video data before facial landmark extraction [abedi2023cross] or to facial landmark data to improve the generalizability of the ST-GCNs, and incorporating hand landmarks and body joints extracted from video into the developed facial-landmarks-based framework.
@@
[…]

### v2

@@ 4.3.3 Results on the Online SE Dataset
[33] 0.7803 0.8764 0.8008
[13] 0.7637 0.8710 0.7980
[44] 0.7475 0.8890 0.7973
[32] 0.8191 0.8926 0.9018
@@
4.3.4 Interpretation of Results
Fig. 1 displays the interpretation of engagement measurements taken using the proposed method by applying Grad-CAM [43] to the ordinal ST-GCN trained on the training set of the EngageNet dataset. For visualization purposes, three exemplary frames (out of 300) from the beginning, middle, and end of three data samples annotated as Not-Engaged in the validation set of the EngageNet dataset are shown in Fig. 1 (a)-(c). The facial landmarks extracted through MediaPipe are overlaid on the frames, where the color map of the facial landmarks, from blue to red, depicts the class activation map values of the last ST-GCN layer in the trained ST-GCN. The facial landmarks associated with the target class of Not-Engaged at certain frames are colored towards red. In Fig. 1 (a), during the beginning and middle of the video, the first two exemplary frames show the student engaged, with lower class activation map values. At the end of the video, when the student is not looking at the camera (computer screen) and is looking elsewhere, landmarks on the iris, eye, and jawline show higher values, resulting in the model classifying the sample as Not-Engaged. In Fig. 1 (b), the student is not paying attention and is playing with their phone. Jawline, eye, iris, and eyebrow facial landmarks with red colors have higher class activation map values, resulting in the model classifying the sample as Not-Engaged. In Fig. 1 (c), throughout the entire video sample, the head pose of the student is normal and perpendicular to the camera. However, the eyes are almost closed, indicating sleepiness and low arousal. This is detected by the higher class activation map values on the eye and iris landmarks, corresponding to the intensity of AU number 45, which indicates how closed the eyes are. This resulted in the model classifying the sample as Not-Engaged.
Figure 1: Interpretation of engagement measurements taken using the proposed method by applying Gradient-weighted Class Activation Mapping (Grad-CAM) to the Spatial-Temporal Graph Convolutional Network (ST-GCN). Please refer to the last paragraph of subsection 4.3.4 for further details.
5 Discussion
@@
Our research led to the development of a novel deep-learning framework for student engagement measurement. In the proposed framework, sequences of facial landmarks are extracted from consecutive video frames and analyzed by ordinal ST-GCNs to make inferences regarding the engagement level of the student in the video. The successful application of our model to the EngageNet [12] and Online SE [42] datasets not only confirms its efficacy but also establishes a new standard in engagement level classification accuracy, outperforming previous methods. As the sole input information to the developed engagement measurement models, the 3D facial landmarks contain information about head pose [17], AUs [18], eye gaze [20], and affect [3, 19], which are the key indicators of behavioral and affective engagement. The relatively lightweight and fast ST-GCN and its integration with real-time MediaPipe [14] make the proposed framework capable of being deployed on virtual learning platforms and measuring engagement in real-time. The proposed method is privacy-preserving and does not require access to personally identifiable raw video data for engagement measurement. In a real-world deployment, the cross-platform MediaPipe solution [14], running on a web, mobile, or desktop application, extracts facial landmarks from video data on users’ local devices. These non-identifiable facial landmarks are then transferred to a cloud, where they are analyzed by ST-GCNs to measure engagement. The interpretability feature of the proposed method, enabled through Grad-CAM, facilitates understanding which facial landmarks, corresponding to behavioral and affective indicators of engagement, contribute to certain levels of engagement. It also helps identify the specific timestamps at which these contributions occur. This provides instructors with additional information to take necessary actions and promote student engagement. A limitation of the proposed method is its reliance on the quality of facial landmarks detected by MediaPipe. In the context of engagement measurement in virtual learning sessions, an occluded or absent face, and consequently non-detected facial landmarks, correspond to lower levels of engagement or disengagement. Our developed ST-GCN was able to correctly classify most samples with occluded or absent faces, i.e., no facial landmarks, as Not-Engaged. To improve the performance of the proposed method, the following direction could be investigated: analyzing facial landmarks with more advanced ST-GCNs, which are equipped with attention mechanisms and trained through contrastive learning techniques and applying augmentation techniques to video data before facial landmark extraction [45] or to facial landmark data to improve the generalizability of ST-GCNs.
@@
This research was funded by the Natural Sciences and Engineering Research Council of Canada.